OpenStudy (glittter22):

A quadratic equation is shown below: 9x2 − 16x + 60 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 4x2 + 8x − 5 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (5 points)

OpenStudy (benlindquist):

dis ill figure out

OpenStudy (glittter22):

thanks for your help

OpenStudy (glittter22):

@benlindquist @hartnn any ideas?

hartnn (hartnn):

the radicand or discriminant for \(ax^2+bx+c =0\) is \(b^2-4ac\) can you find the a,b,c values first?

OpenStudy (glittter22):

a=9 b=-16 and c=60

hartnn (hartnn):

good! so what is \(b^2-4ac = (-16)^2 - 4\times9\times 60 = .. ?\)

OpenStudy (glittter22):

-1904

hartnn (hartnn):

yes. so the discriminant came out to be negative. Do you already know what that means ?

OpenStudy (glittter22):

no..

hartnn (hartnn):

If the discriminant is negative, that means the roots of the quadratic equation are complex conjugate in nature.

hartnn (hartnn):

In other words, both of the roots are NOT real roots.

OpenStudy (glittter22):

so does that mean that the equation has unapplicable answers

hartnn (hartnn):

You can either mention : A. 2 Complex conjugate solutions or B. 2 non-real solutions both will be accepted.

hartnn (hartnn):

it really depends on what (if) you have been already taught by your teacher.

OpenStudy (glittter22):

ok thanks a lot. I already have something for part B but I just want to make sure I got it correct. Give me a minute to type it.

hartnn (hartnn):

sure :)

OpenStudy (inkyvoyd):

yeah, think of it this way - numbers are not actually real. They aren't present in the real world. You can have 2 of something, but not 2 itself. While most would agree there are properties of things that can be modelled with such a "twoness", there you can't have "two". The same applies with complex numbers. The things that complex numbers wind up modelling, however, are different and generally not taught until at least your second year in college. For now, you can simply state that this equation has 2 complex solutions and that no real solutions exist... This is analogous to saying that you can't express a solution in rational numbers (common when you have a square root).

OpenStudy (glittter22):

thx @inkyvoyd

OpenStudy (glittter22):

@hartnn this is part b First we factor 4x^2+8x-5=0 4x^2+10x-2x-5=0 2x(2x+5)-(2x+5)=0 (2x+5)(2x-1)=0 2x+5=0 or 2x-1=0 x= -2.5 x=0.5

hartnn (hartnn):

very good work! :) its correct. \(\huge \checkmark \)

OpenStudy (glittter22):

THank you so much for your help! :)

hartnn (hartnn):

welcome ^_^